OpenStudy (anonymous):

Write an equation of the line that passes through point (5,6) and has a slope of 2.

OpenStudy (anonymous):

y-6=2(x-5) y=2x-4

OpenStudy (anonymous):

\(\ \huge y=mx+b \). \(\ \huge 6=2(5)+b \). \(\ \huge 6=10+b \). \(\ \huge -4=b \). So, \(\ \huge y=2x-4 \).